Use the following information for the Problems below. (The following information applies to the questions displayed below.) T$21 per hr.) 6,237, Factory overhead (297,000 hrs. Total standard cost $16,929,000 Actual costs incurred during the current q. Compute the direct materials cost variance, including its price and quantity variances. Q = Actual Quantity SQ = Standard Q- Compute the direct labor cost variance, including its rate and efficiency variances. H = Actual Hours H= Standard Hours R =Compute the overhead controllable and volume variance Controllable Variance Actual overhead Budgeted overhead Controllable va
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
1474000 $7.40 $10,907,600.00 1474000 $4.40 $6,485,600.00 1485000 $4.40 $6,534,000.00
$4,422,000.00 U $48,400.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $4,422,000.00 U
Direct material quantity variance $48,400.00 F
Direct material cost variance $4,373,600.00 U

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
292000 $12.50 $3,650,000.00 292000 $14.00 $4,088,000.00 297000 $14.00 $4,158,000.00
$438,000.00 F $70,000.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $438,000.00 F
Direct Labor Efficiency variance $70,000.00 F
Direct labor cost variance $508,000.00 F

Solution 3:

Controllable Variance
Actual overhead $5,837,000.00
Budgeted overhead $5,841,000.00
Controllable variance $4,000.00 F
Fixed overhead volume variance
Budgeted fixed overhead $3,168,000.00
Fixed overhead cost applied $3,564,000.00
Fixed overhead volume variance $396,000.00 F